APACHE POINT OBSERVATORY SDSS 2.5M OBSERVING LOG Friday October 30, 2026 (SJD 61344) ---- Observing Team ---- Kaike Pan ---- Observing Plan ---- Fri 30 Oct 2026 61344 0.68 01:08 01:22 12:11 20:57 07:47 --- dark until 04:23 ( 3.0 h), then bright ( 8.0 h, w/ 0.2 h for 12 deg design) ---- Observing Summary ---- The entire night was lost to rain, high humidity, clouds. ---- Observing Log ---- Afternoon: ========= BOSS collimation was checked, no adjustment made.23:16:58 Hartmann collimate output sp1: offset: r= -1649 (Out of focus); b= 1.8 (In Focus) sp1: pred. move: spAverageMove= -1113 sp1: pred. resid: r= -535, b= -1.7, txt= OK, spTemp = 6.8 BOSS cal, APOGEE evening cal. FVC dark, GFA bias, dark are taken, checked. Robots are exercised, confirmed work well. All systems, actors are checked out, confirmed work fine. High humidity, no evening ventilation. Night: ===== Lost to bad weather. ---- Mirror Numbers ---- PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A: -5.4290 -5.4260 Axial B: -5.0390 -5.0390 Axial C: -5.2810 -5.2780 Trans D: -0.8750 -0.8710 Lateral E: 0.3240 0.3245 Lateral F: 0.5335 0.5335 GALILS Commanded: 11450 -1050 5900 26550 -14400 -13850 Actual: 11450 -1050 5900 26550 -14400 -13850 SETMIR VALUES PrimDesOrient: 0.0 -43.05 -8.77 -751.82 -556.2 PrimOrient: -1.06 -43.05 -8.77 -751.82 -556.2 SECONDARY: ---------- Focus: 0.0000 Air temp: 0.00 Altitude: 30.0 MIGS TONIGHT NOMINAL Axial A: -7.9560 -7.9570 Axial B: -6.1120 -6.1140 Axial C: -6.0750 -6.0790 Trans D: 0.2000 0.2050 GALILS Commanded: 363800 387100 345400 -1500 1800 Actual: 321870 350040 299552 -1500 1800 SETMIR VALUES SecDesOrient: 257.0 14.2 5.9 -72.0 -12.76 SecOrient: 256.99 14.21 5.92 -71.78 -12.63 ---- Telescope Status ---- N/A Started ventilation N/A encl25m : open N/A encl25m : closed No interlocks bypassed Status at: 11:50Z Telescope stowed at: 121.0, 30.0, 0.0 mount Instrument mounted: FPS Counterweights at: LN2 autofill systems: Connected and turned on 180L LN2 dewar scales: SP1 249 lbs, 24.8 psi FPS Status ========== CP10 Loop = 7.47 L/min GFA Loop = 2.45 L/min Supply pressure = 43.5 psi Return pressure = 28.2 psi Chiller Status ============== Set temperarature = 8.0 C Flow = 5.0 Gpm ---- Software Used ---- -- STUI software: STUI : 1.7.21 alerts : 2.3.0 boss : 6.0.5a1 -boss.specMec : 3.0.4 apo : 4.1.0 fliswarm : 0.7.0 jaeger : 1.11.2a1 kaiju : 1.4.1 coordio : 1.18.1 fps_calibrati : 2025.12.04+3e7e2d8 cherno : 1.2.2 astrometrynet : 0.97 hal : 1.4.1 hartmann : 2.0.2 hub : 5.0.1a0 mcp : 2.3.1 -mcp.plc : 35 -mcp.azFiduci : v1_154 -mcp.altFiduc : v1_154 -mcp.rotFiduc : v1_154 tcc : 1.4.0 apogee : not availble apogeecal : not availble apogeeql : 2.0.0a0 STUIAdditions : FAILED Boson : 0.7.0 sdss-obstools : 3.9.3 TPM : 1.2-31-gd7eb MCP : 2.3.1 PLC : V35 Fiducials : v1_154 idlspec2d : 6.3.0dev Kronos : 2.1.0a roboscheduler : 2.1.0 ---- Time Tracking ---- *=*=*= Observatory: APO *=*=*= MJD: 61344 Fill night time in Hours below *=*=*= Total bright time : 8.0 *=*=*= Bright time, queued observations : 0 *=*=*= Bright time, commissioning/engineering: 0 *=*=*= Bright time lost to weather : 8.0 *=*=*= Bright time lost to technical issues : 0 *=*=*= Total dark time : 3.0 *=*=*= Dark time, queued observations : 0 *=*=*= Dark time, commissioning/engineering : 0 *=*=*= Dark time lost to weather : 3.0 *=*=*= Dark time lost to technical issues : 0